OpenStudy (anonymous):

Simplify 27 minus 3 times x all over 5 times x minus 45.. A:−1 B:negative 3 over 5 C:negative 24 over 41 D:27 over 45

OpenStudy (anonymous):

\[\frac{27-3x}{5x - 45} = \frac{-3(x - 9)}{5(x - 9)} = \frac{-3}{5}\]